The function f(z)=|z|^{2} is differentiable


1
Expert's answer
2021-08-17T14:19:01-0400

"f(z)=|z|^{2}= z \bar{z}\\"

If f is differentiable at "z=z_0"

Then "\frac{\partial f}{\partial \bar{z}}|_{z=\bar{z}}=0\\"

Now

"\frac{\partial f}{\partial \bar{z}}=z\\
for \space z(\not=0) \in \nsubseteq \frac{\partial f}{\partial \bar{z}} \not=0\\
\implies f(z)=|z|^2" is not differential at "z(\not=)=0 \in \nsubseteq"

To check at z=0

"f'(0)= \lim\limits_{z \to 0 }\frac{f(z)-f(0)}{z-0}\\
f'(0)= \lim\limits_{z \to 0 }\frac{z \bar{z}-0}{z}\\
f'(0)=0"

Hence f(z) is differentiable at z=0 only.